OpenStudy (anonymous):

properties of logarithms: 3^x=5^x+2

OpenStudy (anonymous):

\[3^x=5^{x+2}?\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\ln(3^x)=\ln(5^{x+2})\] \[x \ln 3=(x+2)\ln 5\] \[x \ln 3=x \ln5+2\ln5\] \[x(\ln3-\ln5)=2\ln5\]\[x=\frac{\ln5}{\ln3-\ln5}\]